At a small fair, the first day of ticket sales 600 children’s tickets (x) and 300 adult tickets (y) for a total of $3000. On the second day the fair took in $4250 by selling 900 children’s tickets and 400 adult tickets.
Answer by mananth(16946)

At a small fair, the first day of ticket sales 600 children’s tickets (x) and 300 adult tickets (y) for a total of $3000. On the second day the fair took in $4250 by selling 900 children’s tickets and 400 adult tickets.
Assume you have to find the price of adult ticket and child ticket
At a small fair, the first day of ticket sales 600 children’s tickets (x) and 300 adult tickets (y) for a total of $3000.
600x+300y=3000
divide by 100
6x+3y=30...........................1
On the second day the fair took in $4250 by selling 900 children’s tickets and 400 adult tickets.
900x+400y=4250
divide by 100
9x +4y = 42.5.............................2
Multiply (1) by 4 and (2) by 3
we get
24x+12y =120
27x+12y = 127.50
Subtract
-3x = -7.50
x = 2.50
Plug x value in (1)
6x+3y=30
6*2.50+3y=30
15+3y=30
3y=15
y=5
The price of an adult ticket is $5 and the price of a child's ticket is $2.50.